MY SON HAS THIS PROBLEM TO DO AND I CAN NOT FIGURE IT OUT. CAN YOU PLEASE EXPLAIN HOW TO DO IT. THANKS 3/8=9/X x=?
Answer by wilft1(217)

3/8 = 9/x
your numerator is going from 3 to 9
9 is divisible by 3, so multiply the first denominator by 3 to get 24
3/8 = 9/24
9/24 simplifies back to 3/8